# Reductions in local certification

### v1

@@ 4. Non-$k$-colorability
Given a proof labeling scheme for non-$k$-colorability in $G^{\prime}$, our goal is to assign to each vertex $u\in V(G)$ the list of certificates of a subset $C_{u}$ of vertices of $G^{\prime}$, together with their identifiers, in such a way that $u$ can simulate the verification procedure of a subset $V_{u}$ of vertices of $G^{\prime}$. In this case, our choice for $C_{u}$ and $V_{u}$ is the following: for any $u\in V(G)$, we set $C_{u}=V(G^{\prime})$ (that is, we assign to each vertex of $G$ the list of the certificates of all the vertices of $G^{\prime}$ together with their identifier), and $V_{u}$ consists of $u$ together with the $k-3$ universal vertices of $G^{\prime}$. This allows each vertex $u\in V(G)$ to simulates the verification process of the corresponding vertex in $G^{\prime}$, since its local view in $G^{\prime}$ consists of its neighbors in $G$, together with the universal vertices. Each vertex $u\in V(G)$ also simulates the verification of the $k-3$ universal vertices (whose local view is the whole graph). Now each vertex $u\in V(G)$ accepts the instance if and only if it would have accepted the instance in $G^{\prime}$ and each of the universal vertices would also have accepted the instance in $G^{\prime}$. The size of the certificates is at most $n$ times the size of the certificates for non-$k$-colorability, so by Theorem 3.1 this shows that non-$k$-colorability requires certificates of size $\tfrac{1}{n}\cdot\Omega(n^{2}/\log n)=\Omega(n/\log n)$.
It immediately feels that adding universal vertices is very far from being a local operation, and that this should be avoided as far as possible. Let us now consider a second (more local) way of defining the graph $G^{\prime}$.
@@ 5. Reductions from $\mathcal{P}$ to $\mathcal{P}^{\prime}$
    If there exists a local reduction from $k$-colorability in graphs of maximum degree $k+\lceil\sqrt{k}\rceil-1$ to $\mathcal{P}$ with local expansion $O(n^{\delta})$ for some $10\leqslant\delta<$ 1, and global expansion $O(n^{\gamma})$, then $\overline{\mathcal{P}}$ has local complexity $\Omega\big{(}n^{(1-\delta)/\gamma}/\log n\big{)}$. In particular, since $\gamma\leqslant\delta+1$, $\overline{\mathcal{P}}$ has local complexity $\Omega\big{(}n^{\frac{2}{\delta+1}-1}/\log n\big{)}$.
@@
    By contradiction, assume that there exists a certification scheme for $\overline{\mathcal{P}}$ with certificates of size $o\big{(}n^{(1-\delta)/\gamma}/\log n\big{)}$. By Theorem 5.2, there exists a certification scheme for non-$k$-colorability in graphs of maximum degree $k+\lceil\sqrt{k}\rceil-1$ with certificates of size $o\big{(}n^{\delta}\cdot\frac{(n^{\gamma})^{(1-\delta)/\gamma}}{\log n}\big{)}=o\big{(}\frac{n}{\log n}\big{)}$, which contradicts Theorem 4.1 and Remark 4.2.
@@
Let $G\in\mathcal{C}$ be an $n$-vertex graph in which the vertices are given unique identifiers on $O(\log n)$ bits. First, using a renaming technique from [2], we can assume that the unique identifiers $(\mathrm{id}(u))_{u\in V(G)}$ of $G$ are in $\{1,\ldots,n\}$, at a cost of $O(\log n)$ additional bits per certificates. Since $s(n)=\Omega(\log n)$ and $\beta(n)\geqslant n$, we have $\log n=O\big{(}\alpha(n)\cdot s(\beta(n))\big{)}$, so it does not affect the final bound. We also assume that every vertex knows $n$ (because it can also be certified at a cost of $O(\log n)$ bits in the certificates using a spanning tree, see [7] for more details).
Certification. The certification of $\mathcal{P}$ for $G$ consists in the following. The certificate given by the prover to each vertex consists in a table. For every $u\in V(G)$, we will denote this table by $\mathsf{Table}(u)$. To construct this table, the prover first computes the graph $G^{\prime}=f_{\mathcal{P},\mathcal{P}^{\prime}}(G)$, and for every vertex $u\in V(G)$, it determines the sets $C_{u}$ and $V_{u}$. The entries of $\mathsf{Table}(u)$ are indexed by the identifiers in $G^{\prime}$ of the vertices in $C_{u}$. For every $t\in C_{u}$, we will denote by $\mathsf{Table}(u)[t]$ the entry of $\mathsf{Table}(u)$ indexed by the identifier of $t$. The prover writes in $\mathsf{Table}(u)[t]$ the certificate that $t$ would get in the certification scheme for $\mathcal{P}^{\prime}$ in $G^{\prime}$.
@@
To prove that $G^{\prime}$ satisfies $\mathcal{P}$, it is sufficient to show that there exists an assignment of certificates to the vertices of $G^{\prime}$ such that the verification procedure for $\mathcal{P}^{\prime}$ accepts at every vertex. Let $t\in V(G^{\prime})$. By property Item 3a and Item 3b, there exists $u\in V(G)$ such that $t\in C_{u}$. Let us show that, for every other vertex $w\in V(G)$ such that $t\in C_{w}$, we have $\mathsf{Table}(u)[t]=\mathsf{Table}(w)[t]$. By property Item 3d, the subgraph of $G$ induced by the set $\{v\in V(G)\;|\;t\in C_{v}\}$ is connected. So there exists a path $u_{1},\ldots,u_{k}$ in $G$ such that $u_{1}=u$, $u_{k}=w$, and $t\in C_{u_{i}}$ for every $i\in\{1,\ldots,k\}$. For every $i\in\{1,\ldots,k-1\}$, since $u_{i}$ did not reject at step (ii) of the verification procedure, we have $\mathsf{Table}(u_{i})[t]=\mathsf{Table}(u_{i+1})[t]$. This implies $\mathsf{Table}(u)[t]=\mathsf{Table}(w)[t]$.
@@
All the problems that we study later in this section are mentioned in the book by Garey and Johnson [13], but the reference for each of them is either an unpublished result or a personal communication. Thus, we present here some reductions that we found, which are maybe the same as the ones cited in [13], but to our knowledge, these reductions do not appear in any prior publication. Note that, for the problem of determining whether there exists a cubic subgraph, some reductions were already published (see e.g. [19, 20]), but to our knowledge they are all different from the one we present here (we do not use these previous reductions here, because they are not sufficiently local).
@@ 5.1. Domatic number
The property of having domatic number at most two has local complexity $\Omega(n/\log n)$, even in graphs of maximum degree at most 8.
@@
Let $\mathcal{P}$ be the property of having domatic number at least 3. We will show that there exists a local reduction from 3-colorability in graphs of maximum degree 4 to $\mathcal{P}$ with local expansion $O(1)$, and the result will follow from Corollary 5.3. Let $G$ be a graph of maximum degree at most 4, with $n$ vertices having unique identifiers in $\{1,\ldots,n\}$.
The construction of the graph $G^{\prime}=f_{\text{3-col},\mathcal{P}}(G)$ is the following. For each edge $uv\in E(G)$, we add a new vertex $w$ connected only to $u$ and $v$ (see Figure 6).
$u$$v$$u$$v$$w$
Figure 6. The reduction from 3-colorability to domatic number at least 3.
The identifier of each vertex in $G^{\prime}$ indicates if it is a vertex of $G$ (and which one), or a vertex added for an edge of $G$ (and which edge). This information can be encoded on $O(\log n)$ bits.
For each vertex $u$ of $G$, we define $C_{u}:=V_{u}$ as the set containing $u$ and all the vertices corresponding to an edge $uv$ in $G$ (see Figure 7).
@@
Figure 7. Some graphs $G$ and $G^{\prime}=f_{\text{3-col},\mathcal{P}}(G)$.
Let us show that $G$ admits a proper 3-coloring if and only if the domatic number of $G^{\prime}$ is at least 3. Assume that $G$ is 3-colorable. Then, we create a partition of the vertices of $G^{\prime}$ in the following way: we keep the partition into color classes, and for each new vertex $w$ corresponding to an edge $uv$ of $G$, we assign it to the part to which neither $u$ nor $v$ belongs to. See Figure 8 for an example. Each part of the partition is a dominating set: indeed, each vertex of $G^{\prime}$ belongs to some triangle formed by two original adjacent vertices of $G$ and the new vertex corresponding to the edge, and by construction, each of the three parts of the partition dominates this triangle. Conversely, assume that the vertices of $G^{\prime}$ can be partitioned into at least 3 dominating sets. Then, we obtain a proper 3-coloring of $G$ by restricting this partition to the original vertices of $G$: indeed, let $uv$ be an edge of $G$, and $w$ be the vertex of $G^{\prime}$ corresponding to this edge. Since there are only three vertices in $N[w]$ (which are $u$, $v$ and $w$), these 3 vertices are in different parts of the partition and the partition consists of exactly 3 dominating sets (otherwise, some part does not dominate $w$). In particular, $u$ and $v$ are in different parts, and the corresponding 3-coloring of $G$ is indeed proper.
@@
Figure 8. The transformation from a proper 3-coloring of $G$ to a partition in 3 dominating sets of $G^{\prime}=f_{\text{3-col},\mathcal{P}}(G)$. In the drawing of $G^{\prime}$, the colors correspond to the parts of the partition.
Finally, let us show that the conditions of the definition of local reduction are satisfied. We have already seen that Item 1 and Item 2 both hold. Properties Item 3a, Item 3b, Item 3e and Item 3f simply follow from the definition of $V_{u}$ and $C_{u}$. For Item 3c, since $G$ has maximum degree at most 4, for all $u\in V(G)$ we have $|C_{u}|\leqslant 5$ and thus $|C_{u}|=O(1)$. For Item 3d, if $C_{u}\cap C_{v}\neq\emptyset$, then $u$ and $v$ are neighbors. So all conditions of a local reduction are verified, and Corollary 5.3 gives us a lower bound of order $\Omega(n/\log n)$ on the local complexity of the problem. Since $G$ has maximum degree at most 4, $G^{\prime}=f_{\text{3-col},\mathcal{P}}(G)$ has maximum degree at most 8. So this lower bound holds even for graphs of maximum degree at most 8. ∎
@@ 5.2. Cubic subgraph
Then, we define an tag, represented on Figure 10 (this gadget has already been used in [19] in a different reduction, which was not sufficiently local for our purposes). A tag has only one vertex of degree 1, that we call the root. All the other vertices have degree 3, and we call them the internal vertices.
@@ 5.4. Monochromatic triangles
Let us define a proper 3-coloring of $G$ in the following way. By Claim 5.12, for every $u\in V(G)$, there is a unique $i\in\{1,2,3\}$ such that the main color of $T^{i}_{u}$ is blue. Let $c(u)$ be this unique $i\in\{1,2,3\}$, and let us prove that $c$ is a proper 3-coloring of $G$. By contradiction, assume that there exists $uv\in E(G)$ such that $c(u)=c(v)$. Then, the triangle $T^{c(u)}_{uv}$ is monochromatic (because its edges are colored with the main colors of $T^{c(u)}_{u}$, $T^{c(v)}_{v}$ and $T^{0}_{u}$, respectively, which are all blue), and that is a contradiction. So $G$ is indeed 3-colorable.
@@ 7. Variant of our theorems in a bounded-degree class
In this subsection, we give an example of a lower bound for which the proof is not follow from a direct application of our general results on local reductions, because the definition of local reduction requires a small tweak to be applied efficiently here. This lower bound is stated in the following theorem. We then show that the bound on the maximum degree is close to optimal: if the maximum degree is decreased by two, then the local complexity of the problem drops from $\Omega(n/\log n)$ to $O(\log n)$.
@@
It was proved by Molloy and Reed [16, Theorem 5] that for sufficiently large $k$, a graph $G$ of maximum degree $k+\lceil\sqrt{k}\rceil-3$ is not $k$-colorable if and only if there exists a vertex $v\in V(G)$ such that the subgraph of $G$ induced by $N[v]$ is not $k$-colorable. We use this to produce a proof labeling scheme for non-$k$-colorability as follows. The prover identifies a vertex $v$ such that the subgraph of $G$ induced by $N[v]$ is not $k$-colorable, and each vertex $u$ in the graph receives as a certificate the list $L(u)$ of all neighbors of $u$ (as the graph has bounded maximum degree, this list of identifiers takes $O(\log n)$ bits). Using the lists $\{L(u):uv\in E(G)\}$, the vertex $v$ knows the subgraph of $G$ induced by $N[v]$, and can thus check that this subgraph is indeed not $k$-colorable. ∎

### v2

@@ 4. Non-$k$-colorability
Given a proof labeling scheme for non-$k$-colorability in $G^{\prime}$, our goal is to assign to each vertex $u\in V(G)$ the list of certificates of a subset $C_{u}$ of vertices of $G^{\prime}$, together with their identifiers, in such a way that $u$ can simulate the verification procedure of a subset $V_{u}$ of vertices of $G^{\prime}$. In this case, our choice for $C_{u}$ and $V_{u}$ is the following: for any $u\in V(G)$, we set $C_{u}=V(G^{\prime})$ (that is, we assign to each vertex of $G$ the list of the certificates of all the vertices of $G^{\prime}$ together with their identifier), and $V_{u}$ consists of $u$ together with the $k-3$ universal vertices of $G^{\prime}$. This allows each vertex $u\in V(G)$ to simulate the verification process of the corresponding vertex in $G^{\prime}$, since its local view in $G^{\prime}$ consists of its neighbors in $G$, together with the universal vertices. Each vertex $u\in V(G)$ also simulates the verification of the $k-3$ universal vertices (whose local view is the whole graph). Now each vertex $u\in V(G)$ accepts the instance if and only if it would have accepted the instance in $G^{\prime}$ and each of the universal vertices would also have accepted the instance in $G^{\prime}$. The size of the certificates is at most $n$ times the size of the certificates for non-$k$-colorability, so by Theorem 3.1 this shows that non-$k$-colorability requires certificates of size $\tfrac{1}{n}\cdot\Omega(n^{2}/\log n)=\Omega(n/\log n)$.
The resulting bound of $\Omega(n/\log n)$ is much weaker than the bound of $\Omega(n^{2}/\log n)$ for non-3-colorability. The main reason is that adding universal vertices is very far from being a local operation, and this should be avoided as much as possible. Let us now consider a second (more local) way of defining the graph $G^{\prime}$.
@@ 5. Reductions from $\mathcal{P}$ to $\mathcal{P}^{\prime}$
    If there exists a local reduction from 3-colorability in graphs of maximum degree 4 (or from $k$-colorability in graphs of maximum degree $45k-$ 4) to $\mathcal{P}$ with local expansion $O(n^{\delta})$ for some $10\leqslant\delta<$ 1, and global expansion $O(n^{\gamma})$, then $\overline{\mathcal{P}}$ has local complexity $\Omega\big{(}n^{(1-\delta)/\gamma}/\log n\big{)}$. In particular, since $\gamma\leqslant\delta+1$, $\overline{\mathcal{P}}$ has local complexity $\Omega\big{(}n^{\frac{2}{\delta+1}-1}/\log n\big{)}$.
@@
    By contradiction, assume that there exists a certification scheme for $\overline{\mathcal{P}}$ with certificates of size $o\big{(}n^{(1-\delta)/\gamma}/\log n\big{)}$. By Theorem 5.2, there exists a certification scheme for non-3-colorability in graphs of maximum degree 4, or for non-$k$-colorability in graphs of maximum degree $k+\lceil\sqrt{k}\rceil-1$, with certificates of size $o\big{(}n^{\delta}\cdot\frac{(n^{\gamma})^{(1-\delta)/\gamma}}{\log n}\big{)}=o\big{(}\frac{n}{\log n}\big{)}$. This contradicts Theorem 3.2 and Theorem 4.1, respectively.
@@
Let $G\in\mathcal{C}$ be an $n$-vertex graph in which the vertices are given unique identifiers on $O(\log n)$ bits. First, using a renaming technique from [2], we can assume that the unique identifiers $(\mathrm{id}(u))_{u\in V(G)}$ of $G$ are in $\{1,\ldots,n\}$, at a cost of $O(\log n)$ additional bits per certificates. Since $s(n)=\Omega(\log n)$ and $\beta(n)\geqslant n$, we have $\log n=O\big{(}\alpha(n)\cdot s(\beta(n))\big{)}$, so the renaming does not affect the final bound. We also assume that every vertex knows $n$ (because it can also be certified at a cost of $O(\log n)$ bits in the certificates using a spanning tree, see [8] for more details).
Certification. The certificate given by the prover to each vertex of $G$ consists in a table. For every $u\in V(G)$, this table is denoted by $\mathsf{Table}(u)$. To construct this table, the prover first computes the graph $G^{\prime}=f_{\mathcal{P},\mathcal{P}^{\prime}}(G)$, and for every vertex $u\in V(G)$, it determines the sets $C_{u}$ and $V_{u}$. The entries of $\mathsf{Table}(u)$ are indexed by the identifiers in $G^{\prime}$ of the vertices in $C_{u}$. For every $t\in C_{u}$, we will denote by $\mathsf{Table}(u)[t]$ the entry of $\mathsf{Table}(u)$ indexed by the identifier of $t$. The prover writes in $\mathsf{Table}(u)[t]$ the certificate that $t$ would get in the certification scheme for $\mathcal{P}^{\prime}$ in $G^{\prime}$.
@@
To prove that $G^{\prime}$ satisfies $\mathcal{P}^{\prime}$, it is sufficient to show that there exists an assignment of certificates to the vertices of $G^{\prime}$ such that the verification procedure for $\mathcal{P}^{\prime}$ accepts at every vertex. Let $t\in V(G^{\prime})$. By property Item 3a and Item 3b, there exists $u\in V(G)$ such that $t\in C_{u}$. Let us show that, for every other vertex $w\in V(G)$ such that $t\in C_{w}$, we have $\mathsf{Table}(u)[t]=\mathsf{Table}(w)[t]$. By property Item 3d, the subgraph of $G$ induced by the set $\{v\in V(G)\;|\;t\in C_{v}\}$ is connected. So there exists a path $u_{1},\ldots,u_{k}$ in $G$ such that $u_{1}=u$, $u_{k}=w$, and $t\in C_{u_{i}}$ for every $i\in\{1,\ldots,k\}$. For every $i\in\{1,\ldots,k-1\}$, since $u_{i}$ did not reject at step (ii) of the verification procedure, we have $\mathsf{Table}(u_{i})[t]=\mathsf{Table}(u_{i+1})[t]$. This implies $\mathsf{Table}(u)[t]=\mathsf{Table}(w)[t]$.
@@
All the problems that we study later in this section are mentioned in the book by Garey and Johnson [14], but the reference for each of them is either an unpublished result or a personal communication. Thus, we present here some reductions that we found, which are maybe the same as the ones cited by Garey and Johnson [14], but to our knowledge, these reductions do not appear in any prior publication. Note that, for the problem of determining whether there exists a cubic subgraph, some reductions were already published (see e.g. [22, 23]), but to our knowledge they are all different from the one we present here (we do not use these previous reductions here, because they are not sufficiently local).
@@ 5.1. Domatic number
Let $k\geqslant 2$. The property of having domatic number at most $k$ has local complexity $\Omega(n/\log n)$, even in graphs of maximum degree at most $5k^{2}-4k$.
@@
Let $\mathcal{P}$ be the property of having domatic number at least $k+1$ in graphs of maximum degree $5k^{2}-4k$. We will show that there exists a local reduction from $(k+1)$-colorability in graphs of maximum degree $15k+$ 1 to $\mathcal{P}$ with local expansion $O(1)$, and the result will follow from Corollary 5.3. Let $G$ be a graph of maximum degree at most $15k+$ 1, with $n$ vertices having unique identifiers in $\{1,\ldots,n\}$.
The construction of the graph $G^{\prime}=f_{\text{$(k+1)$-col},\mathcal{P}}(G)$ is the following. For each edge $uv\in E(G)$, we add a clique $K_{uv}$ of size $k-1$, complete to both $u$ and $v$ (see Figure 5 for an example).
$u$$v$$u$$v$$K_{uv}$
Figure 5. The reduction from $(k+1)$-colorability to domatic number at least $k+1$. The clique $K_{uv}$ contains $k-1$ vertices.
The identifier of each vertex in $G^{\prime}$ indicates if it is a vertex of $G$ (and which one), or a vertex in a clique $K_{uv}$ added for an edge $uv$ of $G$ (together with the identifiers of $u$ and $v$, and the numbering of this vertex in $K_{uv}$). This information can be encoded on $O(\log n)$ bits.
For each vertex $u$ of $G$, we define $C_{u}:=V_{u}$ as the set of vertices of $G^{\prime}$ containing $u$ and all the vertices of $K_{uv}$ corresponding to an edge $uv$ in $G$ (see Figure 6 for an example with $k=2$).
@@
Figure 6. Some graphs $G$ and $G^{\prime}=f_{\text{$(k+1)$-col},\mathcal{P}}(G)$, with $k=2$.
Let us show that $G$ admits a proper $(k+1)$-coloring if and only if the domatic number of $G^{\prime}$ is at least $k+1$. Assume that $G$ is $(k+1)$-colorable. Then, we define a ($k+1)$-coloring of the vertices of $G^{\prime}$ in the following way: each vertex of $G$ retains its color in $G^{\prime}$, and for each edge $uv$ in $G$, we color arbitrarily the clique $K_{uv}$ with the $k-1$ colors distinct from that of $u$ and $v$ (each vertex of the clique is assigned a distinct color). See Figure 7 for an example with $k=2$. Note that by construction, each color class is a dominating set in $G^{\prime}$.
Conversely, assume that the vertices of $G^{\prime}$ can be partitioned into $k+1$ dominating sets. Then, we obtain a proper $(k+1)$-coloring of $G$ by considering the restriction of this partition to the original vertices of $G$: indeed, let $uv$ be an edge of $G$, and $w$ be a vertex of $G^{\prime}$ in the $(k-1)$-clique $K_{uv}$. Since there are only $k+1$ vertices in $N[w]$ (which are $u$, $v$ and all the vertices of $K_{uv}$), these $k+1$ vertices must be in different parts of the partition, since otherwise $w$ would not be dominated by one of the $k+1$ dominating sets. In particular, $u$ and $v$ are in different parts, and the corresponding $(k+1)$-coloring of $G$ is indeed proper.
@@
Figure 7. The transformation from a proper $(k+1)$-coloring of $G$ to a partition into $k+1$ dominating sets of $G^{\prime}=f_{(k+1)\text{-col},\mathcal{P}}(G)$, with $k=2$. In the drawing of $G^{\prime}$, the colors correspond to the parts of the partition.
Let us show that the conditions of the definition of local reduction are satisfied. We have already seen that Item 1 and Item 2 both hold. Properties Item 3a, Item 3b, Item 3e and Item 3f simply follow from the definition of $V_{u}$ and $C_{u}$. For Item 3c, since $G$ has maximum degree at most $15k+$ 1, for all $u\in V(G)$ we have $|C_{u}|\leqslant 1+(k-1)(5k+1)=5k^{2}-4k$ and thus $|C_{u}|=O(1)$. For Item 3d, if $C_{u}\cap C_{v}\neq\emptyset$, then $u$ and $v$ are neighbors. So all conditions of a local reduction are verified, and Corollary 5.3 gives us a lower bound of order $\Omega(n/\log n)$ on the local complexity of the problem. Finally, since $G$ has maximum degree at most $15k+$ 1, $G^{\prime}=f_{\text{3-col},\mathcal{P}}(G)$ has maximum degree at most $5k^{2}-4k$. ∎
Remark 5.5.
In Section 7, we will prove with Corollary 7.3 that the maximum degree in the statement of Theorem 5.4 can be decreased to $k^{2}+k\lceil\sqrt{k+1}\rceil$.
@@ 5.2. Cubic subgraph
Then, we define an tag, represented on Figure 9 (this gadget has already been used by Stewart [22] in a different reduction, which was not sufficiently local for our purposes). A tag has only one vertex of degree 1, that we call the root. All the other vertices have degree 3, and we call them the internal vertices.
@@ 5.4. Monochromatic triangles
Let us define a proper 3-coloring of $G$ in the following way. By Claim 5.13, for every $u\in V(G)$, there is a unique $i\in\{1,2,3\}$ such that the main color of $T^{i}_{u}$ is blue. Let $c(u)$ be this unique $i\in\{1,2,3\}$, and let us prove that $c$ is a proper 3-coloring of $G$. By contradiction, assume that there exists $uv\in E(G)$ such that $c(u)=c(v)$. Then, the triangle $T^{c(u)}_{uv}$ is monochromatic (because its edges are colored with the main colors of $T^{c(u)}_{u}$, $T^{c(v)}_{v}$ and $T^{0}_{u}$, respectively, which are all blue), which is a contradiction. So $G$ is indeed 3-colorable.
@@ 7. Variant of our theorems in a bounded-degree class
In this subsection, we give an example of a lower bound for which the proof does not follow from a direct application of our general results on local reductions, because the definition of local reduction requires a small tweak to be applied efficiently here. This lower bound is stated in the following theorem. We then show that the bound on the maximum degree is close to optimal: if the maximum degree is decreased by two, then the local complexity of the problem drops from $\Omega(n/\log n)$ to $O(\log n)$.
@@
As a consequence of Theorem 7.1, we obtain the following improved version of the second item of Corollary 5.3, which implies in particular that the maximum degree in the statement of Theorem 5.4 can be decreased to $k^{2}+k\lceil\sqrt{k+1}\rceil$.
Corollary 7.3.
Let $\mathcal{P}$ be a graph property (on some graph class $\mathcal{C}$) and let $k\geqslant 3$. If there exists a local reduction from $k$-colorability in graphs of maximum degree $k+\lceil\sqrt{k}\rceil-1$ to $\mathcal{P}$ with local expansion $O(n^{\delta})$ for some $10\leqslant\delta<$ 1, and global expansion $O(n^{\gamma})$, then $\overline{\mathcal{P}}$ has local complexity $\Omega\big{(}n^{(1-\delta)/\gamma}/\log n\big{)}$. In particular, since $\gamma\leqslant\delta+1$, $\overline{\mathcal{P}}$ has local complexity $\Omega\big{(}n^{\frac{2}{\delta+1}-1}/\log n\big{)}$.
@@
It was proved by Molloy and Reed [18, Theorem 5] that for sufficiently large $k$, a graph $G$ of maximum degree $k+\lceil\sqrt{k}\rceil-3$ is not $k$-colorable if and only if there exists a vertex $v\in V(G)$ such that the subgraph of $G$ induced by $N[v]$ is not $k$-colorable. We use this to produce a proof labeling scheme for non-$k$-colorability as follows. The prover identifies a vertex $v$ such that the subgraph of $G$ induced by $N[v]$ is not $k$-colorable, and each vertex $u$ in the graph receives as a certificate the list $L(u)$ of all neighbors of $u$ (as the graph has bounded maximum degree, this list of identifiers takes $O(\log n)$ bits). Using the lists $\{L(u):uv\in E(G)\}$, the vertex $v$ knows the subgraph of $G$ induced by $N[v]$, and can thus check that this subgraph is indeed not $k$-colorable. ∎